OpenStudy (anonymous):

How many times does the graph of y=0.1x intersect the graph of y=sin(2x)?

OpenStudy (dumbcow):

the line y=0.1x is in the range of the sin function from -10 to 10 0.1*-10 = -1 0.1*10 = 1 period of sin(2x) is pi 10/pi = 3.18 so approximately 3 cycles on either side of origin intersecting 2 times per cycle Also at x=0, it double counts so take away 1 6*2 -1 = 12-1 = 11 They intersect 11 times verify by graphing